40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 18 18:13:55 2019, Koji, Update, safety, Laser interlock looks OK 
    Reply  Mon Oct 21 11:44:19 2019, gautam, Update, safety, Interlock reconnected to Innolight controller 
Message ID: 14980     Entry time: Mon Oct 21 11:44:19 2019     In reply to: 14978
Author: gautam 
Type: Update 
Category: safety 
Subject: Interlock reconnected to Innolight controller 

We also took this opportunity to re-connect the interlock to the Innolight controller (after it was disconnected for diagnosing the mysterious NPRO self-shutdowns). The diode pump current was dialled down to 0, the interlock wires reconnected, and then the diode current was ramped back up to the nominal 2.1 A. The fan to cool the unit remains mounted in a flaky way as we couldn't locate the frame Chub had made for a more secure mounting solution. 

It seems like the pointing of the beam out of the laser head varies somewhat after the startup - I had to adjust the pointing into the PMC a couple of times by ~1 full turn of the Polaris mount screws, but the IMC has been locked (mostly) for the last ~16 hours.

Quote:

I've checked the state of the laser interlock switch and everything looked normal.

ELOG V3.1.3-